OpenStudy (anonymous):

How do I do this? 4/5k-5=-7+2/5k

OpenStudy (anonymous):

k=-10/2. k=-5

OpenStudy (anonymous):

Is it 4 / ( 5k -5) = - 7 - 2/ (5k)?